Andy reads 56 more pages of his book on Monday than on Tuesday. He reads 125 pages on Monday. How many pages in all does he read on Monday and Tuesday?
Answer by Edwin McCravy(20055)

Andy reads 56 more pages of his book on Monday than on Tuesday.
M = T + 56

He reads 125 pages on Monday.
M = 125

Substitute 125 for M in this equation

  M = T + 56

which gives:

128 = T + 56

Subtract 56 from both sides of the equation:

 128 = T + 56
 -56      -56
-------------
  72 = T + 0
  72 = T

How many pages in all does he read on Monday and Tuesday?
   M + T   =
 125 + 72  =
      197      <--answer

Edwin